# Picogram-Level Nanoplastic Analysis with Nanoelectromechanical System Fourier Transform Infrared Spectroscopy: NEMS-FTIR

### v2

@@ INTRODUCTION
PS | 1452 | 0.50 | 4.25 | 353 |
PP | 1377 | 0.30 | 8.84 | 102 |
PVC | 1427 | 0.49 | 4.17 | 355 |
While NEMS–FTIR can be used to detect very small masses of analytes, very dilute solutions do require pre–concentration. The largest droplet volume that can be drop cast on a NEMS chip is limited to approximately 500 nL. Considering an LOD of 353 pg for PS (see Table 1), the minimum detectable concentration of PS is 0.7 g/mL for a single 500 nL drop of sample. However, in addition to usual pre–concentration techniques such as solid phase extraction, ultrafiltration, and diafiltration, the NEMS–FTIR chip allows new sample pre–concentration possibilities. Multiple droplets can be deposited consecutively on the NEMS chip with intermediate drying steps to increase the analyte mass deposited to the desired level. Sampling via nebulization can also be used to concentrate larger sample volumes onto the NEMS chip. Based on a typical nebulizer flow rate of 20 L/min, a sampling time of 1 hour, and an aerosol capture efficiency of approximately 50 % for 100 nm particles [38], we estimate that concentrations as low as 0.6 ng/mL could be measurable. When dealing with complex matrices containing high concentrations of organic material relative to the target analyte, matrix simplification or removal may be required to remove spectral interferences or avoid overloading the NEMS chip.
@@
The SEM images in Figure 4A, B provide the close–up views of the perforated area of the NEMS chip sampled with PS, PP, and PVC nanoparticles previously mixed in a 1:1:1 mass ratio.
@@
The NEMS–FTIR spectrum of the mixture in Figure 4C shows the characteristic peaks of each polymer component, highlighted by vertical lines. Despite a deposited mass of only 5 ng per component, the peaks corresponding to PS, PP, and PVC are clearly distinguishable. The exact peak positions of the PS, PP, and PVC nanoparticles can be found in the Supporting Information, section A. In environmental contexts, where mixtures of different polymer micro- and nanoparticles are commonly encountered, the broad spectral range of NEMS–FTIR increases the likelihood of detecting distinct peaks without interference, facilitating the identification of individual components when their spectral features are sufficiently resolved.
Tea bag leachates
@@
Figure 5A shows the schematic of the experimental setup for the quantification of nylon tea bag leachates. The empty tea bag was immersed in 200 mL of water preheated at 95C for 10 minutes.
@@
Figure 5B shows micro- and nanoplastic fragments released from the nylon tea bag into the water during standard tea preparation, which were deposited onto the NEMS chip via nanodrop casting. Smaller fragments were also observed on the membranes after sampling the water from the aged tea bags, where a different sample deposition method, based on aerosolization of the liquid, was used (see Figure S4).
Figure 5C shows the NEMS–FTIR spectra of nylon tea bag leachates prepared in triplicates at two aliquot volumes, 100 nL and 500 nL, deposited onto NEMS chips as-is, without prior concentration steps. A comparison with the ATR–FTIR spectrum of 500 nL of the same tea bag leachate reveals that this volume was insufficient for ATR–FTIR detection, as no characteristic peaks were identified. In contrast, the reference ATR–FTIR spectrum of the nylon tea bag itself exhibits well-defined spectral features that correspond closely to those observed in the NEMS–FTIR spectra of the leachates. The high sensitivity of NEMS–FTIR not only enables the use of very small sample volumes but also significantly reduces preparation time, as no concentration is necessary. Volumes of 500 nL of leachate were deposited on the NEMS chip and dried within approximately 30 minutes using the drop casting method (see Sampling methods in Methods). Since this volume failed to produce a detectable signal with ATR–FTIR, a larger sample would be required, which in turn implies longer drying times and more time-consuming sample preparation.
The raw, uncorrected ATR-FTIR spectrum of the tea bag is shown in Figure S5A, alongside the corrected spectrum to illustrate the effects of the applied automatic ATR correction procedures. As can be observed in Figure 5C (see closeup spectrum in Figure S5B), weak diamond phonon absorption bands originating from the ATR crystal [63] appear in the 2500–1800 cm-1 region. These arise due to the refractive index contrast between the sample and diamond ( [66]). Since the refractive index of nylon-based PA is around 1.50–1.53 [67], the difference () is sufficient to make the crystal phonon features visible in the spectrum [63]. Additionally, CO2 absorption peaks are visible within the same region (Figure S5B), contributing to the observed spectral features. The variations in peak position and peak shapes, as well as the presence of additional diamond absorption bands associated with ATR-FTIR spectra, CO2 peaks and potential water vapour peaks, are all absent from the NEMS–FTIR spectra.
Vertical lines, as shown in Figure 5C, highlight characteristic IR peaks corresponding to the chemical structure of PA [68, 69]. The peak at 1720 cm-1 in the NEMS–FTIR spectrum likely originates from imide groups formed during thermal–oxidative degradation of nylon, induced by brewing the tea bag in 95C water [70]. Exact peak positions can be found in the Supporting Information, section B. The same peaks appear in the NEMS–FTIR spectra of water samples in which individual empty nylon tea bags were immersed and subjected to accelerated aging conditions. As aging progressed, the nylon tea bags degraded further, releasing an increasing amount of PA particles, reflected in the peaks’ increasing intensity as shown in the insets (Figure 5D). This observation is in line with previous studies suggesting that prolonged exposure to environmental aging conditions contributes to the release of polymeric particles into the surrounding medium [71].
@@
The NEMS–FTIR spectra collected in both cases allowed for chemical identification of the material from which the tea bags were made and confirmation that this material was leaching into the tea water (Figure 5C, D). Characteristic IR peaks were observed, matching well with the reference ATR–FTIR spectrum of the original tea bag, as well as with the literature [68, 69], confirming the presence of PA, including the amide I (1641 cm-1) and amide II (1553 cm-1) bands. The amide I band primarily arises from C=O stretching vibrations in the amide group, while the amide II band results from a combination of N–H bending and C–N stretching vibrations, which are distinctive features of PA materials. The comparison with online spectral databases, such as Open Specy [72] and Spectragryph [73], identified the released material as nylon–based PA.
@@ CONCLUSIONS
NEMS-FTIR has proven effective for rapidly identifying and quantifying nanoplastics in aqueous dispersions at picogram-level sensitivity. It was demonstrated that drop-cast or nebulized polymer nanoparticles could be identified from polydisperse mixtures and that the measured signals from deposited PS nanoparticles could be used to estimate the mass of PS present on the NEMS chip.
To demonstrate applicability to real-world samples, NEMS-FTIR was used to identify polymer particles released from nylon tea bags during brewing. Leachates were directly analyzed on NEMS chips without prior sample preparation or preconcentration. Compared to ATR-FTIR, NEMS-FTIR offered significantly higher sensitivity, revealing distinct nylon spectral features from as little as 100 nL of sample, whereas ATR-FTIR failed to show peaks even at five times that sample volume. Accelerated aging experiments on nylon tea bags demonstrated the monitoring of the release of polymer particles under prolonged environmental stress. Unlike ATR-FTIR, which requires complex spectral data corrections, careful crystal selection, and suffers from spectral artifacts, NEMS-FTIR simplifies data interpretation through direct absorption measurements. Measurements performed under high vacuum with a nitrogen-purged optical path minimize spectral interference from ambient CO2 and water vapor, and the SiN NEMS chip material can be used as an intrinsic internal standard. NEMS-FTIR spectra are compatible with standard transmission FTIR spectral databases, facilitating the reliable identification of unknown samples.
Taking advantage of the wide availability of FTIR instrumentation, NEMS-FTIR is ideally suited for routine monitoring and widespread adoption. Its non-destructive nature further allows seamless subsequent analysis by complementary techniques such as SEM, EDX, or O-PTIR for detailed sample characterization. The demonstrated performance and ease of integration with existing workflows highlight NEMS-FTIR as a promising tool for environmental monitoring and nanomaterials analysis in general.
@@ METHODS
PS nanoparticles (⌀100 10 nm, Prod. No. 43302), as a 10 % w/v dispersion in water, were obtained from Sigma–Aldrich (MO, USA). PP (⌀54 nm, Prod. No. PP50) and PVC nanoparticles (⌀262 30 nm, Prod. No. PVC250), both as 1 % w/v dispersion in water, were obtained from Lab261 (CA, USA).
@@
All equipment used was first cleaned with acetone (HPLC Plus grade, Prod. No. 650501, Sigma–Aldrich) and isopropyl alcohol (electronic grade, Prod. No. 733458, Sigma–Aldrich), then rinsed five times with distilled deionized water (DDW, 18 Mcm; Millipore, MA, USA) followed by UHPLC–MS grade water (Prod. No. W81, Thermo Fisher Scientific, MA, USA). Finally, all items were allowed to dry under a laminar flow hood (Envirco, NC, USA) to prevent contamination. All samples were prepared in a fume hood (Secuflow, Waldner Holding, Germany) while wearing nitrile gloves (Prod. No. 0439, Helmut Feldtmann GmbH, Germany). Clean nitrile gloves were also used during all experimental procedures.
Serial dilutions and mixture preparation were carried out using UHPLC–MS grade water. A PIPETMAN P1000 micropipette, with matching DIAMOND D1000 TIPACK tips (Prod. No. F144059M and F171500, respectively, Gilson Incorporated, WI, USA), was used for the preparation. All dilutions were stored in hydrolytic class 3 soda-lime glass vials with snap–on lids (Prod. No. LC84.1 and LC87.1, respectively, Carl Roth GmbH, Germany).
All stock and diluted dispersions were first vortexed for 30 seconds using a Mini Vortexer (Heathrow Scientific, IL, USA) to ensure consistent distribution before proceeding with the successive dilution. PS nanoparticle dispersions were prepared at concentrations of 31.25-2000 g/mL, while PP and PVC dispersions ranged from 31.25 to 500 g/mL. The mixed dispersion of all three types of polymer nanoparticles was prepared in a 1:1:1 mass ratio (PS:PP:PVC). Table 2 summarizes all final concentrations used for each single and mixed sample. Blank samples, consisting of the same UHPLC–MS grade water used for dilution procedures, underwent the same preparation process as the experimental samples.
@@
The tea bags used in the experiment were purchased online (Temu.com, China), and the procedure for releasing plastic particles from the tea bags was adapted from the study by Hernandez et al. [52]. To minimize potential contamination, the tea bags, made of nylon according to the manufacturer, were purchased empty, eliminating the need for additional cutting or removing the tea leaves. The thin cotton string that attaches a tiny label to the tea bag was carefully removed with scissors. The tea bags were then rinsed five times with UHPLC–MS grade water to avoid impurities or loose plastic particles and allowed to dry overnight in a laminar flow hood.
To simulate a typical tea brewing process, 200 mL of UHPLC–MS grade water was poured into an Erlenmeyer flask (Duran, DWK Life Sciences GmbH, Germany), which had been pre–cleaned following the same procedure described in the "Preparation of single and mixed nanoplastic dispersions" section. Once the water reached its boiling point, the flask was removed from the hot plate and a single tea bag was immersed. After ten minutes, the tea bag was carefully removed using clean metal tweezers, and the water was allowed to cool overnight. The flask was covered with aluminum foil to prevent contamination during the cooling period. A blank sample was prepared using the same procedure with the same UHPLC–MS grade water but without the tea bag.
@@
The accelerated aging experiment was carried out in a SimTech Feutron Double Climate Chamber (73 x 77 x 102 cm, Feutron Klimasimulation GmbH, Germany) equipped with two Ultra–Vitalux lamps (300 W, Prod. No. 4008321543929, Osram, Germany) placed 30 cm apart (see Figure S6 in the Supporting Information). These lamps emitted broad-spectrum UV light (UVA: 315-400 nm, UVB: 280-315 nm) with an irradiance of approximately 59.1 W/m2, simulating natural solar radiation to induce degradation.
A metal plate, 32 cm below the lamps, held hydrolytic class 1 soda–lime glass vials with screw caps (Prod. No. LC92.1 and LE03.1, respectively, Carl Roth GmbH, Germany) containing the tea bag samples. The vials allowed good UV transmission, particularly in the UVA range, while minimizing light absorption.
The chamber was operated according to a cyclic schedule: 8 hours of UV exposure at 60°C, followed by 4 hours in the dark at 25°C, mimicking natural day and night temperature fluctuations. This regimen, based on ISO 4892-3 [74] and ASTM G154 [75] standards, lasted 25 days, which is equivalent to approximately one year of natural aging [76]. Seven PA tea bags, each in one glass vial with 30 mL of UHPLC–MS water, were artificially aged, and individual water samples were collected after 1, 2, 5, 10, 15, 20, and 25 day(s) to monitor particle release and degradation. Simultaneously, blanks (water without tea bags) were collected with each respective tea bag sample.

### v3

@@ INTRODUCTION
| PS | 1452 | 0.49 | 4.17 | 351 |
| PP | 1377 | 0.29 | 8.67 | 101 |
| PVC | 1427 | 0.48 | 4.09 | 353 |
The NEMS-FTIR spectra of PP and PVC nanoparticles also display peaks which are in agreement with prior studies using FTIR spectroscopy, [79, 80, 76, 81, 82, 78] as well as with the measurements obtained using early IR spectrophotometers employing prisms. [83] A detailed list of all characteristic peaks of PP and PVC nanoparticles, together with their spectral assignments and literature references, is provided in Table S1 of the Supporting Information. NEMS-FTIR does not suffer from the red shift typically associated with measurements performed by ATR-FTIR, as observed in previous studies. [84] In ATR-FTIR, slight shifts in some peak positions are observed due to the interaction of the evanescent wave with the sample surface, which depends on both the refractive index and wavelength. The wavelength-dependent penetration depth inherent to ATR measurements can also cause lower-wavenumber bands to appear more intense compared to transmission (or absorbance) spectra. [54, 55, 56, 57, 58] This type of artifact is absent in NEMS-FTIR spectra. NEMS-FTIR generates spectra comparable to those obtained with transmission-FTIR spectroscopy, allowing seamless integration with existing spectral libraries and chemometric tools for efficient and accurate analysis. Furthermore, since the nanoplastic particles are analyzed under high vacuum, the resulting spectra are free from spectral interferences from water molecules and carbon dioxide (CO2).
Fig. 2D shows the calibration curves for PS, PP, and PVC model nanoparticles. Several characteristic peaks were evaluated for each nanoplastic type to estimate the lowest limits of detection and highest coefficients of determination (R2). Calibration curves were constructed using the characteristic peaks at 1452 for PS (C-H bending vibrations; see inset in Fig. 2C), 1377 for PP (methyl group, CH3, symmetric bending; see inset in Fig. S5), and 1427 for PVC (methylene group, CH2, bending; see inset in Fig. S6) and the results showed good linearity (). LoDs were calculated from the calibration curves using the relation , where is the standard deviation of method blanks () and is the slope of the corresponding calibration curve. The results are summarized in Table 2. PP exhibited the steepest calibration slope, resulting from the strongest IR band used for the calibration curve, and consequently had the lowest LoD. Detection limits are lower than those achieved by Py-GC/MS, with published values ranging from 1 to 10 ng [16, 18, 17] and are comparable to the performance of state-of-the-art TD-PTR-MS techniques with LOD as low as 340 pg. [20]
The variations in detection limits for different nanoplastic particles did not correlate with nanoparticle size. In the measurements conducted on the chips with higher loadings, the observed nanoparticle agglomerations did not result in deviations in linearity.
While NEMS-FTIR can be used to detect very small masses of analytes, very dilute solutions will require pre-concentration. The largest droplet volume that can be drop cast on a NEMS chip is limited to approximately 500 nL. Considering an LOD of 351 pg for PS (see Table 2), the minimum detectable concentration of PS is 0.7 g/mL for a single 500 nL drop of sample. However, in addition to usual pre-concentration techniques such as solid phase extraction, ultrafiltration, and diafiltration, the NEMS-FTIR chip allows new sample pre-concentration possibilities. Multiple droplets can be deposited consecutively on the NEMS chip with intermediate drying steps to increase the analyte mass deposited to the desired level. Sampling via nebulization can also be used to concentrate larger sample volumes onto the NEMS chip. Based on a typical nebulizer flow rate of 20 L/min, a sampling time of 1 hour, and an aerosol capture efficiency of approximately 50 % for 100 nm particles, [43] we estimate that concentrations as low as 0.6 ng/mL could be measurable. For context, Li et al. [85] reported nanoplastic concentrations in tap water, with particle sizes ranging from 58 to 255 nm, varying between 1.67 and 2.08 ng/mL. When dealing with complex matrices containing high concentrations of organic material relative to the target analyte, matrix simplification or removal may be required to remove spectral interferences or avoid overloading the NEMS chip.
@@
The SEM images in Fig. 3A, B provide the close-up views of the perforated area of the NEMS chip sampled with PS, PP, and PVC nanoparticles previously mixed in a 1:1:1 mass ratio. The particles highlighted in these images were identified based on their size, which corresponds to the known size distributions of the individual nanoplastics used: approximately 100 nm for PS, 54 nm for PP, and 262 nm for PVC.
@@
The NEMS-FTIR spectrum of the mixture in Fig. 3C shows the characteristic peaks of each polymer component, highlighted by vertical lines. Despite a deposited mass of only 5 ng per component, the peaks corresponding to PS, PP, and PVC are clearly distinguishable. The exact peak positions of the PS, PP, and PVC nanoparticles can be found in the Supporting Information, in Table S1. In environmental contexts, where mixtures of different polymer micro- and nanoparticles are commonly encountered, the broad spectral range of NEMS-FTIR increases the likelihood of detecting distinct peaks without interference, facilitating the identification of individual components when their spectral features are sufficiently resolved. The obtained spectra correspond to standard FTIR absorption spectra and can be deconvoluted via chemometrics methods.
Nylon teabag leachates
@@
Fig. 4A shows the schematic of the experimental setup for the quantification of nylon teabag leachates. A single empty nylon teabag was immersed in 200 mL of water preheated at 95C for 10 minutes.
@@
Fig. 4B shows micro- and nanoplastic fragments released from the nylon teabag into the water during brewing, which were deposited onto the NEMS chip using piezoelectric nanodroplet dispenser. Smaller fragments were also observed on the membranes after sampling the water from the aged teabags, where a different sample deposition method, based on aerosolization of the liquid, was used (see Fig. S8).
Fig. 4C shows the NEMS-FTIR spectra of nylon teabag leachates (N=3). 100 nL and 500 nL aliquots from the teabag leachates were deposited onto NEMS chips using the piezoelectric nanodroplet dispenser and the drop casting in combination with the pervaporation method, respectively, without prior concentration steps. A 500 nL aliquot of the teabag leachate was also analysed by ATR-FTIR (see Fig. 4C). When compared with the NEMS-FTIR spectrum of the same 500 nL aliquot, the ATR-FTIR signal remains considerably weaker. Only faint indications of the most prominent nylon-based PA bands in the ATR-FTIR spectrum of the teabag leachate become visible upon amplification (), as shown in Fig. S9 in the Supporting Information. In contrast, the reference ATR-FTIR spectrum of the bulk nylon teabag exhibits well-defined spectral features that correspond closely to those observed in the NEMS-FTIR spectra of the leachates, both for the 100 nL and the 500 nL aliquots (see Fig. 4C).
@@
The NEMS-FTIR spectra collected from the teabag leachate allowed for chemical identification of the material from which the teabags were made and confirmation that this material was leaching into the tea water (see Fig. 4C, D). Characteristic IR peaks, indicated by vertical lines in Fig. 4C were observed, matching well with the reference ATR-FTIR spectrum of the original bulk teabag, as well as with the literature, [86, 87, 88, 89, 90, 91, 92] confirming the presence of nylon-based PA, including the amide I (1642 cm-1) and amide II (1553 cm-1) bands. The amide I band primarily arises from C=O stretching vibrations in the amide group, while the amide II band results from a combination of N-H bending and C-N stretching vibrations, which are distinctive features of nylon-based materials [87, 88, 86]. In addition, several other characteristic nylon-related features were also observed in the NEMS-FTIR spectra. These include the CH2 stretching vibrations at 2860 cm-1 (symmetric) and 2930 cm-1 (asymmetric), which are typical of the methylene groups in the nylon structure [87, 88]. Distinctive peaks were detected at 3087 cm-1, associated with N-H stretching overtones, and at 3300 cm-1, corresponding to the N-H stretching vibrations of the amide group. [87, 88, 86, 89, 84] A weak band at 1720 cm-1, that is absent in the ATR-FTIR spectrum of the bulk teabag, appears in the NEMS-FTIR spectra of the leachate. Its presence is consistent with degradation-related functionalities such as imide [89] or carboxylic acid groups [91, 92], which can form during hot-water exposure and indicate the presence of hydrolysed chain ends or unbound low-molecular-weight oligomers. [69, 93, 67] Exact positions and assignments for all observed peaks can be found in Table S1 in the Supporting Information. The comparison with online spectral databases, such as Open Specy [94] and Spectragryph [95], identified the released material as nylon-based PA.
To compare NEMS-FTIR and ATR-FTIR, a correction factor was applied to the ATR-FTIR spectra. The uncorrected ATR-FTIR spectrum of the teabag is shown in Fig. S10A, alongside the corrected spectrum to illustrate the effects of the applied automatic ATR correction procedure. As shown in Fig. 4C, weak diamond phonon absorption bands originating from the ATR crystal [57] appear in the 2500–1800 cm-1 region. A close-up of this spectrum is provided in Fig. 10B. These arise due to the refractive index contrast between the sample and diamond ( [96]). Since the refractive index of nylon is around 1.50–1.53, [97] the difference () is sufficient to make the crystal phonon features visible in the spectrum. [57] Additionally, CO2 absorption peaks are visible within the same region (see Fig. 10B), contributing to the observed spectral features. The variations in peak position and peak shapes, as well as the presence of additional diamond absorption bands associated with ATR-FTIR spectra, CO2 peaks, and potential water vapor peaks, are all absent from the NEMS-FTIR spectra.
Differences in peak ratios were also observed between the ATR-FTIR spectrum of the bulk teabag and the NEMS-FTIR spectra of the 500 nL leachates (see Fig. 4C). In NEMS-FTIR, the amide I peak at 1642 cm-1 is enhanced compared to the ATR-FTIR spectrum. This likely reflects the differences between the samples, as the ATR-FTIR spectra were obtained from the intact bulk polymer, while NEMS-FTIR spectra were obtained from the leachate enriched in oligomeric nylon species released after exposure of the teabag to hot water. [67, 69]
Nylon peaks appear in the NEMS-FTIR spectra of water samples in which individual empty nylon teabags were immersed and subjected to accelerated aging conditions. As aging progressed, the nylon teabags degraded further, releasing an increasing amount of nylon-based PA oligomers and small fragments, reflected in the peaks’ increasing intensity as shown in the insets in Fig. 4D. This observation is in line with previous studies suggesting that prolonged exposure to environmental aging conditions contributes to the release of polymeric particles into the surrounding medium. [98, 99]
The presence of characteristic nylon peaks could also be observed in brewed samples prepared with lemon balm tea leaves without digestion or ultrafiltration, as shown in Fig. 5A. The spectrum of this complex sample (red trace) shows a superposition of the nylon and lemon balm peaks. The nylon teabag spectrum (black trace) can be recovered by subtracting the signal from the pure lemon balm tea leaves (blue trace). Nylon-related peaks are marked with vertical dashed lines. The spectrum of pure lemon balm tea leaves exhibits characteristic features of plant-based polyphenolic extracts, including a dominant band at 1602 cm-1 (C=C stetching), a carbonyl-related band near 1690 cm-1, aromatic ring vibrations in the 1440-1400 cm-1 region, a C-N stretching band near 1270 cm-1, aliphatic CH/CH2 stretching bands around 2900 cm-1, and a broad O-H stretching feature near 3300 cm-1 [100, 101].
The same lemon balm tea extract was brewed with a teabag, oxidized, and ultrafiltrated to remove the organic matrix [62, 71, 66, 65]. As shown in Fig. 5B, the NEMS-FTIR spectra of the resulting fractions (retentate and permeate) showed distinct differences. Although the retentate represents a concentrated fraction (reduced from 200 mL to 20 mL), its spectrum did not show clearly defined nylon peaks; instead, the observed features were largely dominated by signals originating from the lemon balm tea matrix. [100, 101] This is likely due to residual fractions of the tea leaves remaining in the liquid, with particle sizes exceeding the ultrafiltration membrane’s nominal retaining size ( 1 nm). In contrast, the NEMS-FTIR spectrum of the permeate displayed distinct, well-resolved nylon features despite approximately five-fold dilution due to the inclusion of the washing water (180 mL). These results indicate that smaller nylon-derived species, such as dissolved oligomers, passed through the ultrafiltration membrane. The spectrum shows characteristic peaks at 1553, 1642, 2860, 2930, and 3300 cm-1, corresponding to typical vibrational modes of nylon-based PA and its methylene groups [87, 88, 86]. The types and quantities of species released can also vary depending on how the nylon in the teabag was originally manufactured, including differences in polymerisation conditions or additives.
These observations are in agreement with LC-MS/MS analyses, which revealed the presence of nylon oligomers and monomers in teabag leachates. [67, 69] Comparable findings have been observed for PET teabags, with oligomers detected in the leachates. [68] Furthermore, research on other nylon materials has similarly reported the release of nylon oligomers into aqueous media. [93, 69] These results indicate that the concentration of released plastics in complex samples may be underestimated when filtration and dialysis steps are used during the sample preparation if only the retentate is analysed using imaging techniques.
@@ CONCLUSIONS
NEMS-FTIR has proven effective for rapidly identifying and quantifying nanoplastics in aqueous dispersions with picogram-level sensitivity. It was demonstrated that drop-cast or nebulized polymer nanoparticles could be identified from polydisperse mixtures and that the measured signals from deposited PS nanoparticles could be used to estimate the mass of PS present on the NEMS chip.
To demonstrate applicability to real-world samples, NEMS-FTIR was used to identify polymer particles and smaller oligomers released from nylon teabags during brewing. Leachates were directly analyzed on NEMS chips without prior sample preparation or preconcentration. We also show that for relatively simple aqueous matrices, such as tea infusions, identification of nylon leachates could be achieved without digestion, oxidation, ultrafiltration, or concentration steps. Established sample preparation methods can be used to remove concomitant species from more complex environmental samples, such as soil. Compared to ATR-FTIR, NEMS-FTIR offered significantly higher sensitivity, revealing distinct nylon-based PA spectral features from samples as small as 100 nL. ATR-FTIR, even with a five times larger volume, produced only a faint, barely distinguishable signal. Unlike ATR-FTIR, which requires additional spectral data corrections, and careful crystal selection, NEMS-FTIR simplifies data interpretation through direct absorption measurements. Measurements performed under high vacuum with a nitrogen-purged optical path minimize spectral interference from ambient CO2 and water vapor, and the SiN material of the NEMS chip can be used as an intrinsic internal standard to account for chip-to-chip variability. Accelerated aging experiments on nylon teabags demonstrated the monitoring of the release of polymer particles and oligomers under prolonged environmental stress. It was shown that NEMS-FTIR spectra are compatible with standard transmission FTIR spectral databases, facilitating the reliable identification of unknown samples.
Taking advantage of the wide availability of FTIR instrumentation, NEMS-FTIR is ideally suited for routine monitoring and widespread adoption. Its non-destructive nature further allows seamless subsequent analysis by complementary techniques such as SEM, EDX, or O-PTIR for spatial sample characterization. The demonstrated performance and ease of integration with existing workflows highlight NEMS-FTIR as a promising tool for environmental monitoring and nanomaterials analysis in general.
@@ METHODS
PS nanoparticles (⌀100 10 nm, Prod. No. 43302), as a 10 % w/v dispersion in water, were obtained from Sigma-Aldrich (MO, USA). PP (⌀54 nm, Prod. No. PP50) and PVC nanoparticles (⌀262 30 nm, Prod. No. PVC250), both as 1 % w/v dispersion in water, were obtained from Lab261 (CA, USA).
@@
All equipment used was first cleaned with acetone (HPLC Plus grade, Prod. No. 650501, Sigma-Aldrich) and isopropyl alcohol (electronic grade, Prod. No. 733458, Sigma-Aldrich), then rinsed five times with distilled deionized water (DDW, 18 Mcm; Millipore, MA, USA) followed by UHPLC-MS grade water (Prod. No. W81, Thermo Fisher Scientific, MA, USA). Finally, all items were allowed to dry under a laminar flow hood (Envirco, NC, USA) to prevent contamination. All samples were prepared in a fume hood (Secuflow, Waldner Holding, Germany) while wearing nitrile gloves (Prod. No. 0439, Helmut Feldtmann GmbH, Germany). Clean nitrile gloves were also used during all experimental procedures.
Serial dilutions and mixture preparation were carried out using UHPLC-MS grade water. A PIPETMAN P1000 micropipette, with matching DIAMOND D1000 TIPACK tips (Prod. No. F144059M and F171500, respectively, Gilson Incorporated, WI, USA), was used for the preparation. All dilutions were stored in hydrolytic class 3 soda-lime glass vials with snap-on lids (Prod. No. LC84.1 and LC87.1, respectively, Carl Roth GmbH, Germany).
All stock and diluted dispersions were first vortexed for 30 seconds using a Mini Vortexer (Heathrow Scientific, IL, USA) to ensure consistent distribution before proceeding with the successive dilution. PS nanoparticle dispersions were prepared at concentrations of 31.25-2000 g/mL, while PP and PVC dispersions ranged from 31.25 to 500 g/mL. The mixed dispersion of all three types of polymer nanoparticles was prepared in a 1:1:1 mass ratio (PS:PP:PVC). Table 3 summarizes all final concentrations used for each single and mixed sample. Blank samples, consisting of the same UHPLC-MS-grade water used for dilution procedures, underwent the same preparation process as the experimental samples.
@@
Teabags in water
The teabags used in the experiment were purchased online (Temu.com, China), and the procedure for releasing plastic particles from the teabags was adapted from the study by Hernandez et al. [71]
To minimize potential contamination, the teabags, made of nylon according to the manufacturer, were purchased empty, eliminating the need for additional cutting or removing the tea leaves. These empty teabags were first used to establish a controlled reference system, before introducing the more complex matrix containing tea leaves. The thin cotton string that attaches a tiny label to the teabag was carefully removed with scissors. The teabags were then rinsed five times with UHPLC-MS grade water to avoid impurities or loose plastic particles and allowed to dry overnight in a laminar flow hood.
To simulate a typical tea brewing process, 200 mL of UHPLC-MS grade water was poured into an Erlenmeyer flask (Duran, DWK Life Sciences GmbH, Germany), which had been pre-cleaned following the same procedure described in the "Preparation of single and mixed nanoplastic dispersions" section. Once the water reached its boiling point, the flask was removed from the hot plate and a single teabag was immersed. After 10 minutes, the teabag was carefully removed using clean metal tweezers, and the water was allowed to cool overnight. The flask was covered with aluminum foil to prevent contamination during the cooling period. A blank sample was prepared using the same procedure with the same UHPLC-MS grade water but without the teabag.
Teabags and tea
To evaluate the performance of NEMS-FTIR in a more complex matrix containing natural organic matter, fresh lemon balm tea leaves were brewed together with a nylon teabag. The procedure was adapted from Foetisch et al. [62]
First, 200 mL of UHPLC-MS grade water was boiled, after which approximately 1 g of fresh lemon balm leaves was placed in a metal tea sieve (pore size not specified) and immersed in the water at 95C for 5 min. The resulting infusion was divided into two 25 mL aliquots, each of which was reheated to boiling.
A single nylon teabag was added to one of the aliquots and steeped for 5 minutes before being removed with metal tweezers, mimicking a brewing scenario in which both leaves and teabag were present. The second aliquot, containing only tea leaves, served as a control matrix representing organic material derived from the leaves without the contribution of the teabag. After steeping, both aliquots were allowed to cool for several hours while covered with aluminum foil to avoid contamination. Both resulting infusions were then analyzed by NEMS-FTIR post-brewing, and before any organic digestion and ultrafiltration process.
Organic digestion was performed on both samples by six sequential additions of hydrogen peroxide (H2O2; 30%, Prod. No. CP26.1, Carl Roth GmbH, Germany). The amount of H2O2 added was calculated to maintain an overall concentration of approximately 5%, assuming complete consumption of the previously added H2O2.
After digestion, both samples were subjected to ultrafiltration and washing using an Amicon Stirred Cell (200 mL, Prod. No. UFSC20001, Sigma-Aldrich) fitted with a 3 kDa Ultracel membrane (dia. 63.5 mm, Prod. No. PLBC06210, Sigma-Aldrich). The membrane is characterized by a nominal retaining size of approximately 1 nm. [102] Nitrogen gas at 2.3 mbar was used to push the samples through the membranes. The stirred cell reservoir was filled with UHPLC-MS grade water up to 200 mL and filtered until approximately 20 mL remained. Processing approximately 180 mL of liquid required about 2 h. Both the retentate and the permeate were collected for NEMS-FTIR analysis.
@@
The accelerated aging experiment was carried out in a SimTech Feutron Double Climate Chamber (73 x 77 x 102 cm, Feutron Klimasimulation GmbH, Germany) equipped with two Ultra-Vitalux lamps (300 W, Prod. No. 4008321543929, Osram, Germany) placed 30 cm apart (see Fig. S11 in the Supporting Information). These lamps emitted broad-spectrum UV light (UVA: 315-400 nm, UVB: 280-315 nm) with an irradiance of approximately 59.1 W/m2, simulating natural solar radiation to induce degradation.
A metal plate, 32 cm below the lamps, held hydrolytic class 1 soda-lime glass vials with screw caps (Prod. No. LC92.1 and LE03.1, respectively, Carl Roth GmbH, Germany) containing the teabag samples. The vials allowed good UV transmission, particularly in the UVA range, while minimizing light absorption.
The chamber was operated according to a cyclic schedule: 8 hours of UV exposure at 60°C, followed by 4 hours in the dark at 25°C, mimicking natural day and night temperature fluctuations. This regimen, based on ISO 4892-3 [103] and ASTM G154 [104] standards, lasted 25 days, which is equivalent to approximately one year of natural aging. [105] Seven nylon teabags, each in one glass vial with 30 mL of UHPLC-MS water, were artificially aged, and individual water samples were collected after 1, 2, 5, 10, 15, 20, and 25 day(s) to monitor particle release and degradation. Simultaneously, blanks (water without teabags) were collected with each respective teabag sample.